If 150 mL of a sorbitol solution weigh 170 g, what is its specific gravity?

If a liter of a cough syrup weighs 1285 g, what is its specific gravity?

If 500 mL of ferric chloride solution weighs 650 g, what is its specific gravity?

If 380 g of cottonseed oil measures 415 mL, what is its specific gravity?

If 2 fl. oz. of glycerol weighs 74.1 g, what is its specific gravity?

Five pints of diluted hydrochloric acid weighs 2.79 kg. Calculate its specific gravity.

A pycnometer weighs 21.62 g. Filled with water, it weighs 46.71 g; filled with another liquid, it weighs 43.28 g. Calculate the specific gravity of the liquid.

A glass plummet weighs 14.35 g in air, 11.40 g when immersed in water, and 8.95 g when immersed in sulfuric acid. Calculate the specific gravity of the acid.

What is the weight, in kilograms, of 5 liters of sulfuric acid with a specific gravity of 1.84?

What is the volume, in milliliters, of 1 lb of benzyl benzoate having a specific gravity of 1.120?


1
Expert's answer
2021-10-17T11:24:44-0400

Specific gravity = m / V

1. ρ = 170 / 150 = 1.13 g/ml


2. ρ = 1285 / 1000 = 1.285 g/ml


3. ρ = 650 / 500 = 1.3 g/ml


4. ρ = 380 / 415 = 0.91 g/ml


5. 2 fl. oz.= 59.1 ml

ρ = 74.1 / 59.1 = 1.3 g/ml


6. 5 pints = 2.37 liters

ρ = 2.79 / 2.37 = 1.18 g/ml


7. The amount of water contained in the pycnometer is: 46.71 - 21.62 = 25.09 g

The amount of an unknown liquid in the pycnometer is: 43.28 - 21.62 = 21.66 g

The specific gravity of the liquid: 21.66 / 25.09 = 0.86 g/ml


8. Specific gravity is a ratio of the weight of a substance to the weight of an equal volume of a substance chosen as a standard. The standard for liquid and solids is water while for gas is hydrogen.

Amount of displaced acid = 14.35 g — 8.95 g = 5.4 g

Amount of displaced water = 14.35 g — 11.40 g = 2.95 g

Specific gravity of acid = 5.4 g / 2.95 g = 1.8 g/ml


9. m = specific gravity x V

m = 1.84 x 5 = 9.2 kg


10. V = m / Specific gravity

1 lb = 454 g

V = 454 / 1.120 = 405 ml